# Dynamic-SUPERB Phase-2: A Collaboratively Expanding Benchmark for Measuring the Capabilities of Spoken Language Models with 180 Tasks

### v1

@@ Appendix C Task-level evaluation results
| Task | Metric | Whisper- | LTU-AS | SALMONN | SALMONN | Qwen- | Qwen2- | WavLLM | MU- | GAMA | |
| LLaMA | 7B | 13B | Audio | Audio | LLaMA | ||||||
| Vocal Technique Classification | LLM-C | 7.50% | 0.50% | 3.50% | 5.50% | 5.00% | 13.00% | 7.00% | 2.50% | 2.00% | |
| MARBLE Vocal Technique Detection | LLM-C | 8.00% | 1.00% | 11.00% | 6.50% | 6.50% | 15.50% | 8.00% | 8.50% | 2.50% | |
| Lyric Translation | Sacre Bleu | 2.1093 | 1.0357 | 2.8126 | 0.9941 | 3.3975 | 3.9155 | 1.8969 | 0.1593 | 0.0739 | |
| Children Song Transcript | WER | 0.5868 | 1.0268 | 1.0402 | 1.0099 | 1.2867 | 0.9493 | 1.5548 | 1 | 1.0012 | |
| Song Lyric Recognition | MER | 1.0193 | 1.6453 | 3.5636 | 2.6600 | 1.1841 | 1.1562 | 4.8113 | 1.1753 | 1.6852 | |
| Sound Position Prediction | NAR | 81.25% | 100.00% | 100.00% | 100.00% | 100.00% | 25.00% | 93.75% | 100.00% | 100.00% | |
| Angle diff | 1.4006 | - | - | - | - | 1.1063 | 0.8681 | - | - | ||
| Audio Spatial Distance Prediction | NAR | 50.50% | 59.00% | 62.50% | 100.00% | 85.50% | 58.00% | 44.00% | 97.00% | 71.00% | |
| medae | 0.6651 | 0.9330 | 0.4490 | - | 1.5011 | 0.6915 | 0.9011 | 0.9212 | 0.5000 | ||
| How Far Are You | LLM-C | 26.00% | 13.00% | 4.50% | 8.00% | 30.00% | 26.00% | 32.00% | 18.00% | 2.50% | |
| HEAR Vocal Imitation Classification | LLM-C | 15.17% | 5.50% | 1.17% | 1.83% | 18.67% | 11.17% | 17.50% | 3.83% | 5.83% | |
| Scene Fake Detection | LLM-C | 1.98% | 9.41% | 38.12% | 45.54% | 33.17% | 47.03% | 26.73% | 54.46% | 48.51% | |
| Audio Editing Identification | NAR | 84.38% | 87.50% | 62.50% | 75.00% | 75.00% | 68.75% | 96.88% | 65.62% | 37.50% | |
| Acc | 204.3094 | 53.6938 | 165.0617 | 43.1693 | 184.695 | 154.6202 | 100 | 164.0558 | 154.8122 | ||
| Pitch Extraction by Lyrics | NAR | 37.50% | 94.00% | 47.50% | 73.50% | 65.50% | 84.50% | 76.00% | 98.00% | 68.50% | |
| Token Error Rate | 1.4628% | 1.1667% | 2.2621% | 6.3924% | 1.0645% | 1.6596% | 1.3611% | 1% | 1.9518% | ||
| Instrument Pitch Classification | LLM-C | 0.56% | 0.56% | 0.11% | 0.22% | 0.67% | 18.11% | 0.78% | 0.33% | 0.78% | |
| HEAR Music Transcription | NAR | 92.43% | 100.00% | 99.46% | 100.00% | 96.76% | 96.76% | 96.76% | 100.00% | 100.00% | |
| Token Error Rate | 13.0% | - | 36.0% | - | 16.3% | 8.4% | 24.2% | - | - | ||
| HEAR Percussion Instruments Tonic Classification | LLM-C | 2.00% | 2.50% | 12.70% | 13.40% | 11.30% | 11.80% | 12.80% | 12.50% | 5.70% | |
| Audio Deep Fake Detection | LLM-C | 5.56% | 22.06% | 22.06% | 26.19% | 24.44% | 21.43% | 9.21% | 20.95% | 9.84% | |
| Singing Voice Deepfake Detection | LLM-C | 8.25% | 24.56% | 22.79% | 38.31% | 31.43% | 27.90% | 30.84% | 39.10% | 25.54% | |
| Multichannel Sound Event Understanding | LLM-C | 38.73% | 4.23% | 37.32% | 26.06% | 38.03% | 41.55% | 21.83% | 4.93% | 11.27% | |
| Audio Segment Retrieval | NAR | 96.88% | 100.00% | 87.50% | 90.62% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | |
| IoU | 0.0000 | 0.0000 | 0.1238 | 0.0691 | 0.0000 | 0.0000 | 0.0000 | 0.0000 | 0.0000 | ||
| HEAR Sound Event Detection | LLM-C | 0.00% | 0.00% | 21.43% | 35.71% | 7.14% | 14.29% | 21.43% | 21.43% | 7.14% | |
| Singing Automatic MOS Prediction | NAR | 96.88% | 100.00% | 87.50% | 90.62% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | |
| MSE | 12.1700 | 2.3872 | 10.0777 | 11.2308 | 14.8895 | 1.4485 | 11.4210 | 4.5300 | 2.0800 | ||
| KTAU | -0.3333 | 0.0126 | -0.1105 | 0.0166 | 0.0712 | 0.0099 | 0.0103 | 0.1705 | 0.3539 | ||
| LCC | -0.6747 | -0.0706 | -0.1354 | 0.0360 | 0.0824 | 0.0255 | -0.0216 | 0.1866 | 0.3846 | ||
| SRCC | -0.5000 | 0.0183 | -0.1325 | 0.0231 | 0.0872 | 0.0124 | 0.0118 | 0.2281 | 0.4247 | ||
| Emotion Classification in Songs | LLM-C | 10.00% | 0.00% | 1.67% | 3.33% | 13.33% | 8.33% | 8.33% | 1.67% | 0.00% | |
| MARBLE Emotion Detection | LLM-C | 0.10% | 1.50% | 0.00% | 0.00% | 1.30% | 0.50% | 0.10% | 0.00% | 0.10% | |
| Music Beat Tracking (ASAP) | NAR | 54.55% | 74.55% | 51.82% | 63.64% | 62.73% | 40.91% | 47.27% | 80.00% | 84.55% | |
| Miss Time | 135.6077 | 183.6786 | 118.4047 | 77.9687 | 157.5917 | 76.0904 | 111.7759 | 240.0000 | 93.8751 | ||
| Music Beat Tracking (GTZAN-Rhythm) | NAR | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | |
| MSE | - | - | - | - | - | - | - | - | - | ||
| Pitch Extraction by Lyrics | NAR | 37.50% | 94.00% | 47.50% | 73.50% | 65.50% | 84.50% | 76.00% | 98.00% | 68.50% | |
| Token Error Rate | 146.28% | 116.67% | 226.21% | 639.24% | 106.45% | 165.96% | 136.11% | 100% | 195.18% | ||
| Instrument Pitch Classification | LLM-C | 0.56% | 0.56% | 0.11% | 0.22% | 0.67% | 18.11% | 0.78% | 0.33% | 0.78% | |
| HEAR Music Transcription | NAR | 92.43% | 100.00% | 99.46% | 100.00% | 96.76% | 96.76% | 96.76% | 100.00% | 100.00% | |
| HEAR Music Transcription | Token Error Rate | 1.01% | - | 36.0% | - | 8.34% | 1.69% | 6.25% | - | - | |
| HEAR Percussion Instruments Tonic Classification | LLM-C | 2.00% | 2.50% | 12.70% | 13.40% | 11.30% | 11.80% | 12.80% | 12.50% | 5.70% |
| Task | Metric | Whisper- | LTU-AS | SALMONN | SALMONN | Qwen- | Qwen2- | WavLLM | MU- | GAMA | |
| LLaMA | 7B | 13B | Audio | Audio | LLaMA | ||||||
| Environment Recognition | LLM-C | 7.02% | 38.60% | 42.11% | 19.30% | 56.14% | 63.16% | 40.35% | 10.53% | 31.58% | |
| Emergency Traffic Detection | LLM-C | 9.67% | 25.33% | 23.33% | 27.00% | 28.33% | 27.33% | 25.96% | 7.67% | 27.33% | |

### v2

@@ Appendix C Task-level evaluation results
| Task | Metric | Whisper-LLaMA | LTU-AS | SALMONN 7B | SALMONN 13B | Qwen-Audio-Chat | Qwen2-Audio-7B-Inst. | WavLLM | MU-LLaMA | GAMA-IT |
| Chord Classification | LLM-C | 4.00% | 10.00% | 0.50% | 9.00% | 44.00% | 28.50% | 44.50% | 51.00% | 0.00% |
| MARBLE Key Detection (GiantSteps - Key) | LLM-C | 3.00% | 0.50% | 1.50% | 1.00% | 3.00% | 17.00% | 2.00% | 0.50% | 0.00% |
| HEAR Music Transcription (MAESTRO) | NAR | 92.43% | 100.00% | 99.46% | 100.00% | 96.76% | 96.76% | 96.76% | 100.00% | 100.00% |
| HEAR Music Transcription (MAESTRO) | TER | 1.0065 | - | 36.0000 | - | 8.3438 | 1.6875 | 6.25 | - | - |
| HEAR Percussion Instruments Tonic Classification (Mridangam Tonic) | LLM-C | 2.00% | 2.50% | 12.70% | 13.40% | 11.30% | 11.80% | 12.80% | 12.50% | 5.70% |
| Instrument Pitch Classification (Nsynth) | LLM-C | 0.56% | 0.56% | 0.11% | 0.22% | 0.67% | 18.11% | 0.78% | 0.33% | 0.78% |
| Pitch Extraction By Lyrics (CSD) | NAR | 37.50% | 94.00% | 47.50% | 73.50% | 65.50% | 84.50% | 76.00% | 98.00% | 68.50% |
| Pitch Extraction By Lyrics (CSD) | TER | 146.28% | 116.67% | 226.21% | 639.24% | 106.45% | 165.96% | 136.11% | 100.00% | 195.18% |
| Emotion Classificaiton In Songs (EMOTIFY) | LLM-C | 10.00% | 0.00% | 1.67% | 3.33% | 13.33% | 8.33% | 8.33% | 1.67% | 0.00% |
| MARBLE Emotion Detection (MTG) | LLM-C | 0.00% | 1.50% | 0.00% | 0.00% | 2.70% | 0.80% | 0.80% | 0.80% | 0.10% |
| HEAR Music Genre Classification (ISMIR04) | LLM-C | 15.58% | 7.04% | 31.16% | 44.22% | 33.67% | 37.69% | 16.58% | 27.64% | 2.51% |
| MARBLE Genre Classification (MTG) | LLM-C | 0.00% | 1.60% | 0.10% | 0.20% | 3.20% | 0.10% | 0.20% | 0.20% | 0.00% |
| MARBLE Music Tagging (MagnaTagATune) | LLM-C | 5.50% | 1.00% | 3.50% | 9.00% | 4.50% | 1.50% | 1.00% | 3.50% | 7.50% |
| MARBLE Music Tagging (MTG) | LLM-C | 0.10% | 0.30% | 0.00% | 0.00% | 2.90% | 0.30% | 0.10% | 0.50% | 0.10% |
| Music Genre Classification (FMA) | LLM-C | 10.71% | 1.79% | 9.82% | 16.07% | 8.04% | 15.18% | 9.82% | 9.82% | 4.46% |
| HEAR Percussion Instruments Classification (Beijing Opera Percussion) | LLM-C | 3.39% | 4.24% | 10.17% | 20.76% | 16.10% | 27.12% | 24.15% | 22.03% | 11.44% |
| HEAR Percussion Instruments Stroke Classification (Mridangam Stroke) | LLM-C | 0.10% | 3.80% | 7.60% | 4.90% | 11.50% | 10.80% | 13.70% | 10.30% | 0.10% |
| Instrument Classification (Nsynth) | LLM-C | 1.23% | 9.56% | 13.97% | 14.58% | 17.89% | 54.41% | 11.89% | 6.99% | 10.17% |
| Instrument Combination Recognition (OpenMIC-2018) | LLM-C | 4.20% | 13.51% | 17.72% | 24.02% | 25.83% | 24.92% | 20.42% | 19.52% | 8.11% |
| Instrument Source Classification (Nsynth) | LLM-C | 7.22% | 28.33% | 37.33% | 43.56% | 32.67% | 35.22% | 30.89% | 38.00% | 35.89% |
| MARBLE Instrument Classification (MTG) | LLM-C | 0.40% | 2.30% | 0.00% | 0.00% | 6.40% | 1.40% | 0.20% | 0.60% | 0.20% |
| Singing Automatic MOS Prediction (SingMOS) | NAR | 99.50% | 67.72% | 0.33% | 11.15% | 17.30% | 0.00% | 22.96% | 97.34% | 98.34% |
| Singing Automatic MOS Prediction (SingMOS) | MSE | 12.17 | 2.3872 | 10.0777 | 11.2308 | 14.8895 | 1.4485 | 11.421 | 4.53 | 2.08 |
| Singing Automatic MOS Prediction (SingMOS) | KTAU | -0.3333 | 0.0126 | -0.1105 | 0.0166 | 0.0712 | 0.0099 | 0.0103 | 0.1705 | 0.3539 |
| Singing Automatic MOS Prediction (SingMOS) | LCC | -0.6747 | -0.0706 | -0.1354 | 0.036 | 0.0824 | 0.0255 | -0.0216 | 0.1866 | 0.3846 |
| Singing Automatic MOS Prediction (SingMOS) | SRCC | -0.5 | 0.0183 | -0.1325 | 0.0231 | 0.0872 | 0.0124 | 0.0118 | 0.2281 | 0.4247 |
| MARBLE Beat Tracking (ASAP) | NAR | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% |
| MARBLE Beat Tracking (ASAP) | ERR | - | - | - | - | - | - | - | - | - |
| Music Beat Tracking (ASAP) | NAR | 54.55% | 74.55% | 51.82% | 63.64% | 62.73% | 40.91% | 47.27% | 80.00% | 84.55% |
| Music Beat Tracking (ASAP) | Miss Time | 135.6077 | 183.6786 | 118.4047 | 77.9687 | 157.5917 | 76.0904 | 111.7759 | 240 | 93.8751 |
| Audio Editing Identification (People’s Speech) | NAR | 84.38% | 87.50% | 62.50% | 75.00% | 75.00% | 68.75% | 96.88% | 65.62% | 37.50% |
| Audio Editing Identification (People’s Speech) | ACC | 204.3094 | 53.6938 | 165.0617 | 43.1693 | 184.695 | 154.6202 | 100 | 164.0558 | 154.8122 |
| Scene Fake Detection (ASPIRE) | LLM-C | 22.00% | 46.00% | 48.00% | 37.00% | 49.00% | 54.00% | 49.00% | 20.00% | 41.00% |
| Audio Deep Fake Detection (LJSpeech, WaveFake, MUSDB18HQ) | LLM-C | 8.25% | 24.56% | 22.79% | 38.31% | 31.43% | 27.90% | 30.84% | 39.10% | 25.54% |
| Singing Voice Deepfake Detection (CtrSVDD, ACEKiSing, M4Singer) | LLM-C | 5.56% | 22.06% | 22.06% | 26.19% | 24.44% | 21.43% | 9.21% | 20.95% | 9.84% |
| Audio Duration Prediction (NTUML2021) | NAR | 33.00% | 37.00% | 10.50% | 55.00% | 0.00% | 0.00% | 0.50% | 46.50% | 14.00% |
| Audio Duration Prediction (NTUML2021) | MSE | 13.3657 | 39.1111 | 31652.7877 | 2216.7111 | 28.985 | 62.515 | 3581.0352 | 1527.785 | 62.7674 |
| HEAR Music Speech Classification (MAESTRO, Librispeech) | LLM-C | 72.50% | 77.50% | 91.50% | 99.50% | 98.00% | 89.50% | 59.50% | 51.50% | 88.00% |
| Sound Effect Detection (RemFx) | LLM-C | 2.00% | 6.50% | 14.17% | 17.67% | 15.83% | 19.83% | 17.50% | 17.00% | 15.83% |
| Speech Detection (LibriSpeech-TestClean) | LLM-C | 48.00% | 29.00% | 46.00% | 66.00% | 45.50% | 56.50% | 35.00% | 46.50% | 38.50% |
| Speech Detection (LibriSpeech-TestOther) | LLM-C | 52.00% | 31.50% | 46.00% | 66.50% | 44.50% | 55.50% | 37.00% | 45.50% | 36.00% |
| Speech Detection (LJSpeech) | LLM-C | 54.00% | 44.50% | 60.50% | 74.00% | 57.00% | 45.00% | 53.00% | 49.00% | 55.50% |
| Children Song Transcript Verification (CSD) | WER | 58.68% | 102.68% | 104.02% | 100.99% | 128.67% | 94.93% | 155.48% | 100.00% | 100.12% |
| Lyric Translation (SingSet) | Sacre Bleu | 2.1093 | 1.0357 | 2.8126 | 0.9941 | 3.3975 | 3.9155 | 1.8969 | 0.1593 | 0.0739 |
| Song Lyric Recognition (SingSet) | MER | 101.93% | 164.53% | 356.36% | 266.00% | 118.41% | 115.62% | 481.13% | 117.53% | 168.52% |
| MARBLE Vocal Technique Detection (VocalSet) | LLM-C | 8.00% | 1.00% | 11.00% | 6.50% | 6.50% | 15.50% | 8.00% | 8.50% | 2.50% |
| Audio Segment Retrieval (Clotho) | NAR | 80.00% | 78.00% | 2.00% | 4.00% | 12.00% | 4.00% | 14.00% | 82.00% | 6.00% |
| Audio Segment Retrieval (Clotho) | IoU | 8.17% | 22.17% | 8.25% | 8.68% | 4.58% | 7.07% | 6.90% | 0.00% | 10.45% |
| HEAR Sound Event Detection (DCASE2016 Task2) | LLM-C | 0.00% | 0.00% | 21.43% | 35.71% | 7.14% | 14.29% | 21.43% | 21.43% | 7.14% |
| Multi-channel Sound Event Understanding (STARSS23) | LLM-C | 38.73% | 4.23% | 37.32% | 26.06% | 38.03% | 41.55% | 21.83% | 4.93% | 11.27% |
| Animal Classification (WaveSource-Test) | LLM-C | 9.25% | 40.50% | 58.25% | 67.00% | 75.75% | 34.00% | 12.00% | 4.25% | 15.25% |
| Bird Sound Detection (Warblrb10k) | LLM-C | 28.50% | 43.00% | 74.50% | 75.00% | 78.50% | 79.00% | 75.00% | 40.50% | 33.00% |
| Cat Emotion Classification (Cat Sound Classification Dataset V2) | LLM-C | 6.00% | 0.00% | 2.00% | 4.00% | 6.00% | 14.00% | 8.00% | 2.00% | 14.00% |
| Cornell Birdcall Identification | LLM-C | 0.00% | 0.00% | 0.00% | 3.33% | 10.00% | 10.00% | 13.33% | 3.33% | 0.00% |
| Environmental Sound Classification (ESC50-Animals) | LLM-C | 14.50% | 8.00% | 14.00% | 3.00% | 82.00% | 36.00% | 5.00% | 0.00% | 2.00% |
| HEAR Beehive States Classification (Beehive States) | LLM-C | 43.00% | 35.00% | 54.00% | 42.50% | 19.00% | 42.50% | 37.00% | 37.50% | 19.00% |
| Domestic Environment Sound Event Detection (DESED-PublicEval) | NAR | 91.39% | 99.44% | 99.72% | 99.72% | 99.17% | 99.72% | 99.72% | 100.00% | 99.72% |